# Structure Language Models for Protein Conformation Generation

### v1

@@ A.4 Curation of IDP test set
Appendix B Inference of Bidirectional SLMs
B.1 Zero-shot conformation generation of ESM3
The sampling method for the BERT model was first introduced by Wang & Cho (2019), which treats the bidirectional encoder as a Markov random field graphical model and employs Gibbs sampling to generate new sequences. In Hayes et al. (2024), this Gibbs sampling approach is extended to enable simultaneous decoding of multiple positions rather than processing them one at a time. We refer to this latter method as the minibatch Gibbs sampling and take advantages of this sampling strategy for zero-shot conformation generation.
In specific, we leveraged the open-source 1.4B ESM3 protein language model (Hayes et al., 2024) for zero-shot conformation sampling. ESM3 is pre-trained on multi-modal data related to protein and contains several tracks of language model heads for prediction as well as input embeddings, including sequence (amino acid types), structure, secondary structure, SASA and functions. In our main experiment, only the sequence and structure tracks are turned on and the input tokens for other tracks are set to the default value (i.e. [MASK] in each vocabulary). For sampling configurations, we use a sampling schedule of 25 steps and set the sampling temperature to be 1.0 without specially pointing out. During sampling, we also adopted the nucleus (top-p) sampling strategy (Holtzman et al., 2019) to improve the quality of samples with a probability threshold of 0.95. Throughout the experimental Section 5, we use the entropy-ranked minibatch Gibbs sampling to obtain the results for ESM3 (described below).
B.2 Position-Ranked Gibbs Sampling
Unlike one-shot decoding, which maps all masked structure tokens [MASK] to their predictions in a single feed-forward pass, the minibatch Gibbs sampling can adopt various strategies to iteratively select where to unmask at each sampling step from the masked structure tokens of the target protein. These strategies are governed by different ranking functions $f(\cdot)$ applied over the masked positions, as detailed below. The overall inference pipeline (with ranking) is outlined in Algorithm 3. We investigate and compare these ranking strategies in Table B.2, using ESM3 as the base model.

### v2

@@ B.1 Zero-shot conformation generation of ESM3
B.2 Position-Ranked Iterative Decoding
Unlike one-shot decoding, which maps all masked structure tokens [MASK] to their predictions in a single feed-forward pass, the iterative decoding can adopt various strategies to iteratively select where to unmask at each sampling step from the masked structure tokens of the target protein. These strategies are governed by different ranking functions $f(\cdot)$ applied over the masked positions, as detailed below. The overall inference pipeline (with ranking) is outlined in Algorithm 3. We investigate and compare these ranking strategies in Table B.2, using ESM3 as the base model.